# EFC++: Elastic Feature Consolidation with Prototype Re-balancing for Cold Start Exemplar-free Incremental Learning

### v2

@@ VII-B Large-scale Domain- and Class-IL Experiments
Experiments on ImageNet-1K. We compare EFC++ with the nearest competitors in the Cold Start scenario on the large-scale ImageNet-1K dataset (see Table III) using final per-step accuracy. The average incremental accuracy for the Cold Start setting, as well as the per-step and average incremental accuracies for the Warm Start setting, can be found in Appendix F, where findings remain consistent with those on ImageNet-Subset.
@@
Experiments on DN4IL (DomainNet). To further challenge our approach, we consider a domain- and class-incremental setting using the DN4IL dataset, a balanced subset of DomainNet. As shown in Table III, EFC++ improves upon EFC by approximately 1.1% in per-step accuracy and outperforms the remaining state-of-the-art approaches. EFC++ surpasses the best competitor, FeCAM, by roughly 3 points. R-DFCIL struggles here for reasons similar to those observed on ImageNet-1K. Additional comparisons are provided in Appendix F.
@@
VIII Ablations and Additional analysis
@@
VIII-A Ablation Studies
Ablation on EFM regularizer. We evaluate our proposed regularizer, EFM, against three established regularizer: Fisher Information Matrix (E-FIM, Eq. 3), Feature Distillation (FD, Eq. 5), and Knowledge Distillation (KD) [53]. To carry out this comparison, we directly replaced EFM regularization in our training framework without modifying any other components. Specifically, we maintained the same prototype updates and classifier rebalancing procedures. In order to perform these experiments, we carefully validate the hyperparameters associated with each regularizer (see Appendix H for more details on the validation). For these experiments, in addition to the per-step accuracy, we report the average forgetting and the average plasticity (see Eq. 23), to measure the stability and the plasticity induced by each regularizer. As shown in Table IV, EFM achieves the best trade-off between stability and plasticity, leading to the highest final accuracy in both the 10-step and 20-step settings on CIFAR-100. Figure 8 further highlights EFM effectiveness: its average forgetting (left) remains comparable to FD during the incremental steps, while its plasticity (right) is significantly improved. In contrast, while KD and E-FIM exhibit strong plasticity (right), they suffer from severe forgetting (left) after the first task, ultimately degrading their final performance as shown in Table IV.
TABLE IV: Ablation study on regularizers. We report the final average forgetting ($F^{K}$), average plasticity ($\text{PL}^{K}$), and per-step accuracy ($A_{\text{step}}^{K}$) for CIFAR-100 at 10 and 20 steps.
@@ VIII-A Ablation Studies
TABLE V: Ablation study on prototype update. The performance of EFC++ improves with prototype updates in both Warm and Cold Start.
@@
Ablation on prototype update. In Table V we evaluate the effect of prototype updates using Eq. 14 and Eq. 15. Prototype updates significantly improve performance in both Warm Start and Cold Start across most benchmarks, except for ImageNet-Subset WS, where no significant improvement is observed.
VIII-B Analysis of Feature Drift in Warm and Cold Start
@@
VIII-C EFM Spectral Analysis and Hyperparameters
@@
Moreover, the horizontal line compares the magnitude of the eigenvalues across tasks with the damping term. Recalling that the results reported for EFC++ were obtained using parameter values of $\lambda_{\text{EFM}}=10$ and $\eta=0.1$, we observe that the plasticity constraints $\lambda_{\text{EFM}}\nu_{i}>\eta$ (described in Section III-C) are satisfied for nearly every $\nu_{i}>0$, clearly demonstrating that our regularizer effectively constrains the features in a non-isotropic manner, distinguishing it from feature distillation.
To provide empirical proof of the importance of calibrating such hyperparameters (i.e., avoiding too low a value of $\eta$ or degenerating into feature distillation), we also conducted a grid search testing 25 different combinations of $\lambda_{\text{EFM}}$ and $\eta$. Figure 11 (left) shows that the selected combination achieves the peak value in terms of final per-step accuracy. In particular, it reveals that excessively low values of $\eta$ lead to weaker regularization and increased forgetting (middle), whereas excessively high values significantly reduce plasticity (right). See the Appendix G for a more comprehensive analysis on other scenarios.
VIII-D Training Time and Memory Requirements
@@
Memory requirements. EWC and LwF do not use prototypes, so no additional storage is required. FeTrIL, PASS, and SSRE need to store only class means. R-DFCIL increases storage requirements by including a generative deep inversion model with about 33M parameters (roughly 125.89 MB at 32-bit precision). EFC, EFC++, and FeCaM must store both class means and covariances. Comparing their memory requirements to R-DFCIL’s, storing 100 class covariance matrices of size 512×512 (as in CIFAR-100 with a ResNet-18 backbone) requires about 100 MB – slightly less than R-DFCIL. However, as the number of classes increases, the storage requirements for EFC, EFC++, and FeCaM grow with respect to R-DFCIL. In our previous work [29], we showed that these costs can be mitigated through covariance approximation, thus eliminating the need to store a covariance matrix per class—or by applying a low-rank approximation. This significantly reduces storage without significant performance degradation.
VIII-E Analysis of the limitations of EFC++
TABLE VI: Impact of Mean and Covariance Drift on EFC++ (CIFAR-100 Cold Start).
@@ VIII-E Analysis of the limitations of EFC++
Since EFC++ does not update class covariances across incremental learning steps, we aim to understand how covariance drift affects the final performance. Table VI compares different update strategies of class statistics: our approach (class means estimated via EFM and fixed covariances, highlighted in cyan) against an idealized setting using real mean and real covariances, computed from old task data fed in the current task backbone. While this strategy is infeasible in incremental learning – since past data is unavailable – it serves as an upper bound for performance and helps quantify the limitations of our prototype rehearsal strategy.
The results suggest that updating the means via EFM and using real class covariances (third row) does not provide significant benefit. However, replacing our estimated class means with real ones, while keeping the covariance fixed, leads to a substantial improvement: 2.92% and 5.33% with respect to having fixed statistics in 10 and 20 steps (fourth row). This confirm that challenges highlighted in Section VIII-B, that our approach encounters when estimating the real class drift in Cold Start. Using real class covariances significantly improves performance only as the number of steps increases (fifth row), but this improvement is much smaller compared to the impact of using real class means. In Appendix I we show that similar conclusion holds for Tiny-ImageNet and ImageNet-Subset.
@@ IX Conclusions
Supplementary Material
Simone Magistri, Tomaso Trinci, Albin Soutif–Cormerais, Joost van de Weijer, Andrew D. Bagdanov
@@ Appendix A Scope and Summary Notation
  • $\mathcal{P}_{1:t-1}$: prototypes accumulated up to the task $t-1$.
@@ Appendix B The Empirical Feature Matrix
In this appendix we derive an analytic formulation of the local features matrix defined in the main paper in Section III-C. To simplify the notation, we recall its definition without explicitly specifying the task, as it is unnecessary for the derivation:
@@
$=I_{m}-\mathbbm{1}_{m}\cdot\begin{bmatrix}\sigma_{1}(z)\\ \vdots\\ \sigma_{m}(z)\\ \end{bmatrix}^{\top}=I_{m}-P,$ (27)
where $I_{m}\in\mathbb{R}^{m\times m}$ is the identity matrix and $\mathbbm{1}_{m}$ represent the vector with all entries equal to $1$.
@@
$=WJ(z),$ (29)
where $z=Wf$ for $W$ the classifier weight matrix mapping features $f$ to logits $z$. Combining this with Eq. 27 we have:
@@ Appendix D Dataset, Implementation and Hyperparameter Settings
In this section we expand the discussion of Section VI-A and Section VI-D of the main paper by providing additional information about the dataset and the implementation of the approaches we tested.
D-A Datasets
We perform our experimental evaluation on four standard datasets. CIFAR-100 [65] consists of 60,000 images divided into 100 classes, with 600 images per class (500 for training and 100 for testing). ImageNet-1K is the original ImageNet dataset [67], consisting of about 1.3 million images divided into 1,000 classes. Tiny-ImageNet [66] consists of 100,000 images divided into 200 classes, which are taken from ImageNet and resized to $64\times 64$ pixels. ImageNet-Subset is a subset of the original ImageNet dataset that consists of 100 classes. The images are resized to $224\times 224$ pixels. DNIL [32] is a balance subset of DomainNet and it consists of six domains, each with 100 classes, and a total of 85,000 images. The images are resized to $224\times 224$ pixels. Each domain (i.e. clipart, infograph, painting, quickdraw, real, and sketch) is treated as a different task, resulting in a 6-step scenario used to evaluate the performance when large domain shifts occur.
D-B Implementation and Hyperparameters
For all the methods, we use the standard ResNet-18 [1] backbone trained from scratch.
First Task Optimization Details. We use the same optimization settings for both the Warm Start and Cold Start scenarios. We train the models on CIFAR-100 and on Tiny-ImageNet for 100 epochs with Adam [71] using an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs (as done in [21, 23]). For ImageNet-Subset, we followed the implementation of PASS [21], fixing the number of epochs at 160, and used Stochastic Gradient Descent with an initial learning rate of 0.1, momentum of 0.9, and weight decay of $5e^{\text{-}4}$. The learning rate was reduced by a factor of 0.1 after 80, 120, and 150 epochs. We applied the same label and data augmentation (random crops and flips) for all the evaluated datasets. For the first task of each dataset, we use self-rotation as performed by [21, 23]. For both ImageNet-1k and DN4IL, we apply the same hyperparameters and training protocol used for the ImageNet-Subset. The batch size is set to 64 for DN4IL and 256 for ImageNet-1k. For ImageNet-1k, self-rotation is deactivated for the first task.
@@ D-B Implementation and Hyperparameters
  • EWC [10]: We used the implementation of [14]. Specifically, we configured the coefficient associated to the regularizer as $\lambda_{\text{E-FIM}}=5000$ and the fusion of the old and new importance weights is done with $\alpha=0.5$. For the incremental steps we fix the total number of epochs to $100$ and we use Adam optimizer with an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs.
@@
  • LwF [53]: We used the implementation of [14]. In particular, we set the temperature parameter $T=2$ as proposed in the original work and the parameter associated to the regularizer $\lambda_{\text{LwF}}$ to $10$. For the incremental steps we fix the total number of epochs to $100$ and we use Adam optimizer with an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs.
@@
  • PASS [21]: We follow the implementation provided by the authors. It is an approach relying upon feature distillation and prototypes generation. Following the original paper we set $\lambda_{\text{FD}}=10$ and $\lambda_{\text{pr}}=10$. In the original code, we find a temperature parameter, denoted as $T$, applied to the classification loss, which we set to $T=1$. As provided in the original paper, for the incremental steps we fix the total number of epochs to 100 and we use Adam optimizer with an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs
@@
  • SSRE [22]: We follow the implementation provided by the authors. It is an approach relying upon feature distillation and prototypes generation. Following the original paper, we set $\lambda_{\text{FD}}=10$ and $\lambda_{\text{pr}}=10$. In the original code, we find a temperature parameter, denoted as $T$, applied to the classification loss, which we set to $T=1$. Following the code, for the incremental steps we fixed the total number of epochs to $60$ and used Adam with an initial learning rate of $2e^{\text{-}4}$ and fixed weight decay of $5e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 epochs.
@@
  • FeTrIL [26]: We follow the official code provided by the authors. During the incremental steps, it uses a Linear SVM classifier working on the pseudo-features extracted from the frozen backbone after the first task. We set the SVM regularization $C=1$ and the tolerance to $0.0001$ as provided by the authors.
@@
  • FeCAM [27]: We follow the implementation provided by the authors. During the steps, it uses Mahalanobis-based distance for classification on the features extracted from the frozen backbone after the first task and transformed via Tukey’s Ladder of Powers transformation. To improve and stabilize the Mahalanobis-based distance classification, it applies covariance normalization and covariance shrinkage. Following the original implementation, we set the hyperparameter of Tukey’s Ladder of Powers transformation, $\lambda$, to $0.5$ and the hyperparameters of the covariance shrinkage, $\gamma_{1}$ and $\gamma_{2}$, to $1$.
@@
  • Deep Inversion Methods: We trained ABD [31] and R-DFCIL [30] following their official implementations. The only difference from the original papers is that CIFAR-100 and Tiny-ImageNet were tested using a ResNet-32 [1] backbone. However, for a fair comparison, we replaced it with ResNet-18. Regarding optimization details, we follow the same settings originally used by the authors for ImageNet-Subset, where they employed ResNet-18. Specifically, we train all benchmarks for 100 epochs for ABD and 120 epochs for R-DFCIL, using SGD with an initial learning rate of $0.1$ for CIFAR-100, Tiny-ImageNet, and DN4IL, and $0.005$ for ImageNet-Subset and ImageNet-1K, following the original learning rate schedule [^1] We chose not to test ABD on DN4IL and ImageNet-1K because it generally performs worse than R-DFCIL on all other datasets.. We apply a fixed weight decay of $1e^{\text{-}4}$ and a momentum of $0.9$. The batch size is set to $128$ for CIFAR-100 and Tiny-ImageNet, while for DN4IL, ImageNet-Subset, and ImageNet-1K, it is $64$. We use the same deep inversion network, running 10,000 inversion iterations for CIFAR-100 and Tiny-ImageNet, and 50,000 iterations for ImageNet-Subset and ImageNet-1K.

### v3

@@ 7.2 Large-scale Domain- and Class-IL Experiments
Experiments on ImageNet-1K. We compare EFC++ with the nearest competitors in the Cold Start scenario on the large-scale ImageNet-1K dataset (see Table 3) using final per-step accuracy. The average incremental accuracy for the Cold Start setting, as well as the per-step and average incremental accuracies for the Warm Start setting, can be found in Appendix F, where findings remain consistent with those on ImageNet-Subset.
@@
Experiments on DN4IL (DomainNet). To further challenge our approach, we consider a domain- and class-incremental setting using the DN4IL dataset, a balanced subset of DomainNet. As shown in Table 3, EFC++ improves upon EFC by approximately 1.1% in per-step accuracy and outperforms the remaining state-of-the-art approaches. EFC++ surpasses the best competitor, FeCAM, by roughly 3 points. R-DFCIL struggles here for reasons similar to those observed on ImageNet-1K. Additional comparisons are provided in Appendix F.
@@
8 Ablations and Additional analysis
@@
8.1 Ablation Studies
Ablation on EFM regularizer. We evaluate our proposed regularizer, EFM, against three established regularizer: Fisher Information Matrix (E-FIM, Eq. 3), Feature Distillation (FD, Eq. 5), and Knowledge Distillation (KD) lwf . To carry out this comparison, we directly replaced EFM regularization in our training framework without modifying any other components. Specifically, we maintained the same prototype updates and classifier rebalancing procedures. In order to perform these experiments, we carefully validate the hyperparameters associated with each regularizer (see Appendix H for more details on the validation). For these experiments, in addition to the per-step accuracy, we report the average forgetting and the average plasticity (see Eq. 23), to measure the stability and the plasticity induced by each regularizer. As shown in Table 4, EFM achieves the best trade-off between stability and plasticity, leading to the highest final accuracy in both the 10-step and 20-step settings on CIFAR-100. Figure 8 further highlights EFM effectiveness: its average forgetting (left) remains comparable to FD during the incremental steps, while its plasticity (right) is significantly improved. In contrast, while KD and E-FIM exhibit strong plasticity (right), they suffer from severe forgetting (left) after the first task, ultimately degrading their final performance as shown in Table 4.
Table 4: Ablation study on regularizers. We report the final average forgetting ($F^{K}$), average plasticity ($\text{PL}^{K}$), and per-step accuracy ($A_{\text{step}}^{K}$) for CIFAR-100 at 10 and 20 steps.
@@ 8.1 Ablation Studies
Table 5: Ablation study on prototype update. The performance of EFC++ improves with prototype updates in both Warm and Cold Start.
@@
Ablation on prototype update. In Table 5 we evaluate the effect of prototype updates using Eq. 14 and Eq. 15. Prototype updates significantly improve performance in both Warm Start and Cold Start across most benchmarks, except for ImageNet-Subset WS, where no significant improvement is observed.
8.2 Analysis of Feature Drift in Warm and Cold Start
@@
8.3 EFM Spectral Analysis and Hyperparameters
@@
Moreover, the horizontal line compares the magnitude of the eigenvalues across tasks with the damping term. Recalling that the results reported for EFC++ were obtained using parameter values of $\lambda_{\text{EFM}}=10$ and $\eta=0.1$, we observe that the plasticity constraints $\lambda_{\text{EFM}}\nu_{i}>\eta$ (described in Section 3.3) are satisfied for nearly every $\nu_{i}>0$, clearly demonstrating that our regularizer effectively constrains the features in a non-isotropic manner, distinguishing it from feature distillation.
To provide empirical proof of the importance of calibrating such hyperparameters (i.e., avoiding too low a value of $\eta$ or degenerating into feature distillation), we also conducted a grid search testing 25 different combinations of $\lambda_{\text{EFM}}$ and $\eta$. Figure 11 (left) shows that the selected combination achieves the peak value in terms of final per-step accuracy. In particular, it reveals that excessively low values of $\eta$ lead to weaker regularization and increased forgetting (middle), whereas excessively high values significantly reduce plasticity (right). See Appendix G for a more comprehensive analysis on other scenarios.
8.4 Training Time and Memory Requirements
@@
Memory requirements. EWC and LwF do not use prototypes, so no additional storage is required. FeTrIL, PASS, and SSRE need to store only class means. R-DFCIL increases storage requirements by including a generative deep inversion model with about 33M parameters (roughly 125.89 MB at 32-bit precision). EFC, EFC++, and FeCaM must store both class means and covariances. Comparing their memory requirements to R-DFCIL’s, storing 100 class covariance matrices of size 512×512 (as in CIFAR-100 with a ResNet-18 backbone) requires about 100 MB – slightly less than R-DFCIL. However, as the number of classes increases, the storage requirements for EFC, EFC++, and FeCaM grow with respect to R-DFCIL. In our previous work magistri2024elastic , we showed that these costs can be mitigated through covariance approximation, thus eliminating the need to store a covariance matrix per class—or by applying a low-rank approximation. This significantly reduces storage without significant performance degradation.
8.5 Analysis of the limitations of EFC++
Table 6: Impact of Mean and Covariance Drift on EFC++ (CIFAR-100 Cold Start).
@@ 8.5 Analysis of the limitations of EFC++
Since EFC++ does not update class covariances across incremental learning steps, we aim to understand how covariance drift affects the final performance. Table 6 compares different update strategies of class statistics: our approach (class means estimated via EFM and fixed covariances, highlighted in cyan) against an idealized setting using real mean and real covariances, computed from old task data fed in the current task backbone. While this strategy is infeasible in incremental learning – since past data is unavailable – it serves as an upper bound for performance and helps quantify the limitations of our prototype rehearsal strategy.
The results suggest that updating the means via EFM and using real class covariances (third row) does not provide significant benefit. However, replacing our estimated class means with real ones, while keeping the covariance fixed, leads to a substantial improvement: 2.92% and 5.33% with respect to having fixed statistics in 10 and 20 steps (fourth row). This confirms that challenges highlighted in Section 8.2, that our approach encounters when estimating the real class drift in Cold Start. Using real class covariances significantly improves performance only as the number of steps increases (fifth row), but this improvement is much smaller compared to the impact of using real class means. In Appendix I we show that similar conclusion holds for Tiny-ImageNet and ImageNet-Subset.
@@ 9 Conclusions
Acknowledgements.
This work was supported by funding by the European Commission Horizon 2020 grant #951911 (AI4Media), TED2021-132513B-I00 and PID2022-143257NB-I00 funded by MCIN/AEI/ 10.13039/501100011033, by the Italian national Cluster Project CTN01_00034_23154 “Social Museum Smart Tourism”, by the European Union project NextGenerationEU/PRTR, and by the Spanish Development Fund FEDER.
Data Availability  All data employed in this paper are publicly available. In our public repository (https://github.com/simomagi/elastic_feature_consolidation), we provide detailed instructions to download and use the data for the experiments.
Appendix
@@ Appendix A Scope and Summary Notation
  • $\mathcal{P}_{1:t-1}$: prototypes accumulated up to the task t1t-1.
@@ Appendix B The Empirical Feature Matrix
In this appendix we derive an analytic formulation of the local features matrix defined in the main paper in Section 3.3. To simplify the notation, we recall its definition without explicitly specifying the task, as it is unnecessary for the derivation:
@@
$=I_{m}-\mathds{1}_{m}\cdot\begin{bmatrix}\sigma_{1}(z)\\ \vdots\\ \sigma_{m}(z)\\ \end{bmatrix}^{\top}=I_{m}-P,$
where $I_{m}\in\mathbb{R}^{m\times m}$ is the identity matrix and $\mathds{1}_{m}$ represent the vector with all entries equal to $1$.
@@
$=WJ(z),$
where $z=Wf$ for $W$ the classifier weight matrix mapping features $f$ to logits $z$. Combining this with Eq. LABEL:eqn:simple-jacobian we have:
@@ Appendix D Dataset, Implementation and Hyperparameter Settings
In this section we expand the discussion of Section 6 of the main paper by providing additional information about the dataset and the implementation of the approaches we tested.
D.1 Datasets
We perform our experimental evaluation on four standard datasets. CIFAR-100 CIFAR100 consists of 60,000 images divided into 100 classes, with 600 images per class (500 for training and 100 for testing). ImageNet-1K is the original ImageNet dataset ImageNet , consisting of about 1.3 million images divided into 1,000 classes. Tiny-ImageNet TinyImageNet consists of 100,000 images divided into 200 classes, which are taken from ImageNet and resized to $64\times 64$ pixels. ImageNet-Subset is a subset of the original ImageNet dataset that consists of 100 classes. The images are resized to $224\times 224$ pixels. DNIL gowda2023dual is a balance subset of DomainNet and it consists of six domains, each with 100 classes, and a total of 85,000 images. The images are resized to $224\times 224$ pixels. Each domain (i.e. clipart, infograph, painting, quickdraw, real, and sketch) is treated as a different task, resulting in a 6-step scenario used to evaluate the performance when large domain shifts occur.
D.2 Implementation and Hyperparameters
For all the methods, we use the standard ResNet-18 resnet backbone trained from scratch.
First Task Optimization Details. We use the same optimization settings for both the Warm Start and Cold Start scenarios. We train the models on CIFAR-100 and on Tiny-ImageNet for 100 epochs with Adam adam using an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs (as done in pass ; evanescent ). For ImageNet-Subset, we followed the implementation of PASS pass , fixing the number of epochs at 160, and used Stochastic Gradient Descent with an initial learning rate of 0.1, momentum of 0.9, and weight decay of $5e^{\text{-}4}$. The learning rate was reduced by a factor of 0.1 after 80, 120, and 150 epochs. We applied the same label and data augmentation (random crops and flips) for all the evaluated datasets. For the first task of each dataset, we use self-rotation as performed by pass ; evanescent . For both ImageNet-1k and DN4IL, we apply the same hyperparameters and training protocol used for the ImageNet-Subset. The batch size is set to 64 for DN4IL and 256 for ImageNet-1k. For ImageNet-1k, self-rotation is deactivated for the first task.
@@ D.2 Implementation and Hyperparameters
  • EWC ewc : We used the implementation of facil . Specifically, we configured the coefficient associated to the regularizer as $\lambda_{\text{E-FIM}}=5000$ and the fusion of the old and new importance weights is done with $\alpha=0.5$. For the incremental steps we fix the total number of epochs to $100$ and we use Adam optimizer with an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs.
@@
  • LwF lwf : We used the implementation of facil . In particular, we set the temperature parameter $T=2$ as proposed in the original work and the parameter associated to the regularizer $\lambda_{\text{LwF}}$ to $10$. For the incremental steps we fix the total number of epochs to $100$ and we use Adam optimizer with an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs.
@@
  • PASS pass : We follow the implementation provided by the authors. It is an approach relying upon feature distillation and prototypes generation. Following the original paper we set $\lambda_{\text{FD}}=10$ and $\lambda_{\text{pr}}=10$. In the original code, we find a temperature parameter, denoted as $T$, applied to the classification loss, which we set to $T=1$. As provided in the original paper, for the incremental steps we fix the total number of epochs to 100 and we use Adam optimizer with an initial learning rate of $1e^{\text{-}3}$ and fixed weight decay of $2e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 and 90 epochs
@@
  • SSRE ssre : We follow the implementation provided by the authors. It is an approach relying upon feature distillation and prototypes generation. Following the original paper, we set $\lambda_{\text{FD}}=10$ and $\lambda_{\text{pr}}=10$. In the original code, we find a temperature parameter, denoted as $T$, applied to the classification loss, which we set to $T=1$. Following the code, for the incremental steps we fixed the total number of epochs to $60$ and used Adam with an initial learning rate of $2e^{\text{-}4}$ and fixed weight decay of $5e^{\text{-}4}$. The learning rate is reduced by a factor of 0.1 after 45 epochs.
@@
  • FeTrIL fetril : We follow the official code provided by the authors. During the incremental steps, it uses a Linear SVM classifier working on the pseudo-features extracted from the frozen backbone after the first task. We set the SVM regularization $C=1$ and the tolerance to $0.0001$ as provided by the authors.
@@
  • FeCAM goswami2023fecam : We follow the implementation provided by the authors. During the steps, it uses Mahalanobis-based distance for classification on the features extracted from the frozen backbone after the first task and transformed via Tukey’s Ladder of Powers transformation. To improve and stabilize the Mahalanobis-based distance classification, it applies covariance normalization and covariance shrinkage. Following the original implementation, we set the hyperparameter of Tukey’s Ladder of Powers transformation, $\lambda$, to $0.5$ and the hyperparameters of the covariance shrinkage, $\gamma_{1}$ and $\gamma_{2}$, to $1$.
@@
  • Deep Inversion Methods: We trained ABD smith2021always and R-DFCIL gao2022r following their official implementations. The only difference from the original papers is that CIFAR-100 and Tiny-ImageNet were tested using a ResNet-32 resnet backbone. However, for a fair comparison, we replaced it with ResNet-18. Regarding optimization details, we follow the same settings originally used by the authors for ImageNet-Subset, where they employed ResNet-18. Specifically, we train all benchmarks for 100 epochs for ABD and 120 epochs for R-DFCIL, using SGD with an initial learning rate of $0.1$ for CIFAR-100, Tiny-ImageNet, and DN4IL, and $0.005$ for ImageNet-Subset and ImageNet-1K, following the original learning rate schedule [^1] We chose not to test ABD on DN4IL and ImageNet-1K because it generally performs worse than R-DFCIL on all other datasets.. We apply a fixed weight decay of $1e^{\text{-}4}$ and a momentum of $0.9$. The batch size is set to $128$ for CIFAR-100 and Tiny-ImageNet, while for DN4IL, ImageNet-Subset, and ImageNet-1K, it is $64$. We use the same deep inversion network, running 10,000 inversion iterations for CIFAR-100 and Tiny-ImageNet, and 50,000 iterations for ImageNet-Subset and ImageNet-1K.